# Leveraging the true depth of LLMs

### v1

@@ 4 Efficient Parallelization of Blocks
Figure 5: Layer Parallel (LP) implementation of attention. In this diagram, the stacked layers represent different GPUs, the colors indicate the intermediate tensors of different layers and the arrows express linear projections. In this case, the number of GPUs and the number of parallelized layers coincides and is 2, which is the set-up that we use for all our experiments in this work. We see how tensors for the two layers are distributed across the two GPUs. We also notice the reduction step happening at the end, summing the outputs from each GPUs, which differs from the regular computational graph obtained when running layers in parallel as in equation (PAR).
Naive block parallelization. Pure parallelization of two transformer blocks would consist in implementing equation (PAR). To optimize our parallel implementation, we need to leverage efficient GPU kernels by e.g. concatenating together matrices from different blocks. As an example, let’s consider parallelizing ${\bm{y}}_{1}=W_{1}{\bm{x}}$ and ${\bm{y}}_{2}=W_{2}{\bm{x}}$, ${\bm{x}}\in\mathbb{R}^{d_{x}},{\bm{y}}\in\mathbb{R}^{d_{y}},W_{1},W_{2}\in\mathbb{R}^{d_{y}\times d_{x}}$. We can concatenate $W=[W_{1},W_{2}]\in\mathbb{R}^{2*d_{y}\times d_{x}}$ and get $[{\bm{y}}_{1},{\bm{y}}_{2}]=W{\bm{x}}$. However, this would not work if, instead of a shared input ${\bm{x}}$, we had two separate inputs ${\bm{x}}_{1}$ and ${\bm{x}}_{2}$. This is precisely the situation we are in when we apply different layer norms to the input of the two attention blocks. The Feed-Forward Networks (FFN) also have different inputs, in addition to having different layer norms. Those considerations directed us to modify the computational graph of the parallel processing of blocks. While we differ from equation (PAR), we show that our approach nonetheless—and quite surprisingly—works well on already trained models, circumventing the need to train from scratch.
The hardware limits of parallelisms. Another difficulty we faced when parallelizing transformer blocks is the saturation of GPU resources. LLM inference typically saturates GPU resources due to its intensive compute and memory bandwidth requirements. When all the GPU cores are already being utilized by one matrix multiplication, running another matrix multiplication in parallel can be as slow as running them sequentially. This is the case with large transformer models, as even processing a single sequence can fully utilize a GPU’s capabilities. As such, simply attempting to run two blocks in parallel would result in sequential execution, as the scheduler would allocate operations from both blocks to the same job stream. To achieve true parallel execution of layers, we decided to leverage Tensor Parallelism by distributing layer weights across multiple GPUs.
We extend the tensor parallelism scheme introduced in Megatron (Shoeybi et al., 2020) to incorporate our novel Layer Parallelism. Below, we detail our approach for each major component of the transformer block, initially setting aside Layer Normalization considerations and tackling the Multi-Head Attentions (MHA) and the FFN. Our approach is also illustrated in Fig. 2. It is important to note that the resulting computational graph is not numerically equivalent to the original architecture. This numerical discrepancy stems from the positioning of the pre-normalization operations that precede each transformer sub-block.
@@
Layer Parallel Multi-Head Attention. Traditional tensor parallelism in MHA distributes attention heads evenly across GPUs, performing self-attention and output projection locally before gathering results through worker summation. Each GPU processes tensors of dimensions $Q,K,V,att\in\mathbb{R}^{T\times\frac{D}{g}}$, where $T$ is sequence length, $D$ is feature dimension, and $g$ is the number of parallel workers. The local output projection produces $o_{i}\in\mathbb{R}^{T\times D}$ for each worker $i$, with rank $\frac{D}{g}$ until the gather operation restores full rank. To implement Layer Parallelism, we increase the depth of the query, key, and value weight matrices ($W_{Q},W_{K},W_{V}\in\mathbb{R}^{(g_{n}\cdot h_{d})\times D}$) and widen the output projection ($W_{O}\in\mathbb{R}^{D\times(n_{h}\cdot h_{d})}$), where $n_{h}$ represents heads per GPU and $h_{d}$ is head dimensionality. The reduction operation now will simultaneously compute the full-rank output projections and the sum of all parallel layers (Fig. 5(b)).
@@
Layer Parallel Feed Forward Network. Standard tensor parallelism for single-hidden-layer FFNs splits the first layer’s output across devices, generates partial outputs from the second layer, and combines them through reduction. To parallelize two FFN layers, we double the first layer’s output dimensionality and perform separate output projections for each layer. A single reduction operation then serves the dual purpose of computing full outputs for each layer and combining their results, as shown in Fig.  2(b). In summary, Layer Parallelism for FFN just concatenates the up-projection weights and continues as normal TP, allowing for multiple GPUs to be allocated per parallelized layer.
Handling Layer Normalization. Layer Normalization presents unique challenges since these layers were trained on specific input distributions. For MHA pre-normalization, we apply separate normalization on each device. For FFN pre-normalization, we found that linearly interpolating the weights of both FFN pre-norms yielded better perplexity than maintaining separate normalizations. This improvement may stem from the fact that, given we reduce the MHA outputs over the parallelized layers, interpolation effectively combines the expected input distributions of both layers.
@@ 5 Experiments & Results
Effective Depth Par. Layers MMLU PiQA Arc Easy Arc Challenge Winogrande openbooqka hellaswag
Llama 2 7B
32 (Baseline) - 0.4583 0.8009 0.8106 0.5196 0.7411 0.4520 0.7821
27 (Ours) 18-28 0.4625 0.7933 0.8005 0.5094 0.7348 0.4600 0.7782
26 (Ours) 16-28 0.4588 0.7927 0.7976 0.4983 0.7340 0.4460 0.7745
25 (Ours) 14-28 0.4532 0.7851 0.7917 0.4949 0.7340 0.4440 0.7673
24 (Ours) 12-28 0.4083 0.7845 0.7841 0.4839 0.7190 0.4360 0.7578
23 (Ours) 10-28 0.3519 0.7829 0.7677 0.4488 0.6922 0.4240 0.7368
Llama 3.2 3B
28 (Baseline) - 0.5610 0.7992 0.7807 0.4872 0.7214 0.4520 0.7557
24 (Ours) 17-25 0.5508 0.7856 0.7521 0.4753 0.7167 0.4420 0.7384
23 (Ours) 15-25 0.5481 0.7748 0.7399 0.4735 0.7119 0.4200 0.7303
22 (Ours) 13-25 0.4693 0.7666 0.7264 0.4497 0.6914 0.4180 0.7193
21 (Ours) 11-25 0.3890 0.7519 0.6839 0.4061 0.6638 0.4020 0.6847
20 (Ours) 9-25 0.3107 0.7416 0.6481 0.3652 0.6227 0.3620 0.6407
Table 1: 5-shot In-Context Learning accuracies across standard benchmarks. Effective Depth shows the minimum number of sequential operations from input to output. Par. Layers indicates the range of consecutive layers where pairs were processed in with Layer Parallelism.
Finetuning Steps MMLU Rel. MMLU (%)
0 (Baseline) 0.5610 100
0 (13-25 LP) 0.4693 83.6
4096 (Ours) 0.4979 88.8
8192 (Ours) 0.5295 94.4
16384 (Ours) 0.5222 93.1
32736 (Ours) 0.5266 93.8
Table 2: Recovery of MMLU accuracy through finetuning on Llama 3.2 3B with Layer Parallelism applied to layers 13-25. Relative MMLU shows performance as a percentage of the baseline model’s accuracy. The recovered MMLU saturates after fine-tuning for 8k steps.
@@
Figure 6: Perplexity when running pairs of consecutive layers in parallel. Perplexity of Llama2 7B and Llama3.2 3B models on the test set of RedPajama(Together Computer, 2023) when applying Layer Parallelism to $\Delta$ consecutive layers. The parallelized interval for each data point is $[\text{end index}-\Delta,\text{end index}[$.
Figure 7: Wall clock time to complete the following inference tasks: KV Cache pre-filling for a given sequence length, autoregressive generation up to the indicated sequence length, and single token generation with a pre-filled KV Cache of the indicated sequence length. The baseline is the original model with all layers making use of Tensor Parallelism. The Parallel Layers number ($\Delta$) indicates how many layers have been merged using Layer Parallelism (e.g. a $\Delta$ of 4 indicates that 2 groups of 2 layers have been converted to 2 effective layers). The gains in inference speed are directly proportional to the grade of Layer Parallelism. The 1-token generation task for Llama 3.2 3B does not saturate the GPU compute until a sequence length of 2048. Even in this regime, Layer Parallelism benefits from considerable speed-ups.
@@
Experimental protocol. For all our experiments, we use a node with two A100 SXM4 80Gb GPUs, four AMD EPYC 7742 CPUs, and 512Gb of RAM. We test for varying sequence lengths, up to $4096$ (Llama’s context window), with a batch size of $1$ unless indicated otherwise. We consider two models of the Llama family: Llama2 7B, and Llama3.2 3B. We always apply Layer Parallelism of $2$ (one layer to each GPU) on the merged sequential layers and apply Tensor Parallelism as described in (Shoeybi et al., 2020) for the rest. The baselines are fully Tensor Parallel Llama models. For evaluation, we measure the ICL 5-shot accuracies using the lm-eval package (Gao et al., 2024). We test the ICL accuracy of the models on several tasks: MMLU (Hendrycks et al., 2021), PiQA (Bisk et al., 2019), ARC Easy, ARC Challenge, Winogrande (Sakaguchi et al., 2021), OpenBookQA (Mihaylov et al., 2018) and Hellaswag (Zellers et al., 2019). The perplexity (PPL) of the models is always evaluated against a subset of the test set of RedPajama (Together Computer, 2023).
Impact of layer-parallelism on PPL and ICL accuracies. We begin by exploring the evolution of the perplexity when applying layer parallelism to stretches of layers of varying lengths, and starting at different depths. Results in Fig. 6 show how both models—Llama2 7B and Llama3.2 3B—exhibit a common sequence ending index for which the perplexity is minimized, which is 28 and 25 for Llama2 7B and Llama3.2 3B, respectively. Taking this into consideration, in Table 1 we evaluate the In-Context Learning capabilities of models of different effective depths, for which parallelized sequences end at those indices. For Llama 2 7B we observe that applying Layer Parallelism to a sequence greater than 14 layers results in a steep loss of In-Context Learning capabilities in the more challenging benchmarks, like MMLU. Likewise, parallelizing above 10 layers of Llama 3.2 3B sees an even more rapid decrease in performance on difficult benchmarks. The effective depth of both models at the parallel configurations before those sudden drops in performance is 25 and 23, a reduction of 21% and 18% of their original depths, respectively.
@@
Fine-tuning for performance recovery. While Layer Parallelism offers significant speed improvements, the architectural modifications can impact model performance. To address this, we investigated whether fine-tuning could recover the original model’s capabilities. Using Llama 3.2 3B with Layer Parallelism applied to layers 13-25 ($\Delta=12$), we fine-tuned only the parallelized layers on random samples from RedPajama’s training set (Together Computer, 2023). With a batch size of 2 and a learning rate of $41e-$ 4, we observed substantial recovery of model performance. As shown in Table 2, fine-tuning for 8,192 steps improved MMLU accuracy from 83.6% to 94.4% of the baseline performance, demonstrating that much of the model’s original capability can be recovered while maintaining the speed benefits of Layer Parallelism.
@@ 6 Limitations
While our approach demonstrates significant improvements in inference efficiency, several important limitations should be considered.
The effectiveness of our approach exhibits notable variations across model scales. Smaller models show reduced benefits, likely due to their less sparse activation patterns and more tightly coupled layer dependencies. Even in successful cases, we observe a consistent, albeit small, performance degradation compared to the baseline. This degradation becomes more pronounced as model depth increases, suggesting a practical upper limit to the number of layer pairs that can be effectively parallelized.
@@
Determining the ’true’ effective depth—the optimal configuration of parallel layer pairs—remains an open challenge as there is no theoretical framework for predicting the optimal grouping strategy. These limitations highlight important directions for future research, particularly in developing more robust methods for determining optimal layer groupings and investigating the interplay between our approach and other efficiency-oriented techniques.

### v2

@@ 4 Efficient Parallelization of Blocks
Figure 5: LP attention implementation. This diagram shows the implementation of the LP attention from Fig. 2b. The stacked layers represent different GPUs, the colors indicate different layers and the arrows express linear projections. In this case, the number of GPUs and the number of parallelized layers coincides and is two, which is the set-up that we use for all our experiments in this work.
If we naively tried to run two attention or MLP blocks by fusing their execution on a single GPU, we would not see any noticeable improvement on the inference speed. This is due to the intensive compute and memory bandwidth requirements of LLMs. For this reason we focus our attention on the tensor parallel setting, where each module’s weights are split over two or more GPUs. Rearranging the computational graph of two contiguous layers like in Fig. 2b effectively reduces the number of inter-GPU synchronizations by half. Now, each divergent path is computed in parallel over multiple accelerators, and only the single intermediate and final results need to be synchronized. While this approach is not numerically equivalent to (PAR), we nonetheless—and quite surprisingly— show that it works well on already trained models, circumventing the need to train from scratch.
@@
LP Multi-Head Attention. Traditional tensor parallelism in MHA distributes attention heads evenly across GPUs[37], performing self-attention and output projection locally before gathering results through an all-reduce summation. Each GPU processes tensors of dimensions $Q,K,V,att\in\mathbb{R}^{T\times\frac{D}{g}}$, where $T$ is sequence length, $D$ is feature dimension, and $g$ is the number of parallel workers. The local output projection produces $o_{i}\in\mathbb{R}^{T\times D}$ for each worker $i$, with rank $\frac{D}{g}$, and then the all-reduce operation restores full rank. To implement LP, we increase the depth of the query, key, and value weight matrices ($W_{Q},W_{K},W_{V}\in\mathbb{R}^{(g_{n}\cdot h_{d})\times D}$) and widen the output projection ($W_{O}\in\mathbb{R}^{D\times(n_{h}\cdot h_{d})}$), where $n_{h}$ represents heads per GPU and $h_{d}$ is head dimensionality. The reduction operation now will simultaneously compute the full-rank output projections and the sum of all parallel layers (Fig. 5).
@@
LP FFN. Standard tensor parallelism for single-hidden-layer FFNs splits the first layer’s output across devices, generates low-rank outputs from the second layer, and sums them through reduction. To parallelize two FFN layers, we double the first layer’s output dimensionality and perform separate output (low-rank) projections for each layer. A single reduction operation then serves the dual purpose of computing full outputs for each layer and combining their results, as shown in Fig.  2(b). In summary, LP for FFN just concatenates the up-projection weights and continues as normal TP, allowing for multiple GPUs to be allocated per parallelized layer.
Handling LN. Since we asume at least one GPU per parallelized layer, we can assign each original LN to the divergent path that contains the attention and FFN blocks from its original layer.
@@ 5 Experiments & Results
Figure 6: Perplexity when running pairs of consecutive layers in parallel. Perplexity of Llama2 7B and Llama3.2 3B models on the test set of RedPajama[1] when applying Layer Parallelism to $\Delta$ consecutive layers. The parallelized interval for each data point is [end indexΔ,end $[\text{end index}-\Delta,\text{end index}]$.
Table 1: 5-shot In-Context Learning accuracies across standard benchmarks. Effective Depth shows the minimum number of sequential operations from input to output. Par. Layers indicates the range of consecutive layers where pairs were processed in with Layer Parallelism. *ifeval was evaluated on 0-shot performance.
Eff. Depth MMLU PiQA Arc E. Arc C. WinoG OBQA hswag GSM8K ifeval* Avg.
Llama 2 7B
32 (Base) 0.4583 0.8009 0.8106 0.5196 0.7411 0.4520 0.7821 0.6521
27 (Ours) 0.4625 0.7933 0.8005 0.5094 0.7348 0.4600 0.7782 0.6484
26 (Ours) 0.4588 0.7927 0.7976 0.4983 0.7340 0.4460 0.7745 0.6431
25 (Ours) 0.4532 0.7851 0.7917 0.4949 0.7340 0.4440 0.7673 0.6386
24 (Ours) 0.4083 0.7845 0.7841 0.4839 0.7190 0.4360 0.7578 0.6248
23 (Ours) 0.3519 0.7829 0.7677 0.4488 0.6922 0.4240 0.7368 0.6006
Llama 3.2 3B
28 (Base) 0.5610 0.7992 0.7807 0.4872 0.7214 0.4520 0.7557 0.6510
24 (Ours) 0.5508 0.7856 0.7521 0.4753 0.7167 0.4420 0.7384 0.6373
23 (Ours) 0.5481 0.7748 0.7399 0.4735 0.7119 0.4200 0.7303 0.6284
22 (Ours) 0.4693 0.7666 0.7264 0.4497 0.6914 0.4180 0.7193 0.6058
21 (Ours) 0.3890 0.7519 0.6839 0.4061 0.6638 0.4020 0.6847 0.5690
20 (Ours) 0.3107 0.7416 0.6481 0.3652 0.6227 0.3620 0.6407 0.5273
Llama 2 7B (Chat)
32 (Base) 0.4727 0.7769 0.7963 0.4906 0.7206 0.3300 0.5887 0.2297 0.4365 0.5380
27 (Ours) 0.4746 0.7709 0.7803 0.4881 0.7182 0.3380 0.5839 0.2221 0.4736 0.5389
26 (Ours) 0.4724 0.7666 0.7748 0.4701 0.7151 0.3400 0.5779 0.1903 0.4856 0.5325
25 (Ours) 0.4767 0.7633 0.7719 0.4650 0.7009 0.3320 0.5724 0.1478 0.4796 0.5233
24 (Ours) 0.4547 0.7655 0.7567 0.4369 0.6788 0.3040 0.5582 0.0963 0.4149 0.4962
23 (Ours) 0.4271 0.7524 0.7428 0.4113 0.6654 0.3080 0.5406 0.0697 0.3765 0.4771
Llama 3.2 3B (Instruct)
28 (Base) 0.5956 0.7709 0.7929 0.4650 0.7024 0.3080 0.5253 0.6475 0.6715 0.6088
24 (Ours) 0.5914 0.7622 0.7660 0.4488 0.7009 0.3000 0.5131 0.4587 0.6031 0.5716
23 (Ours) 0.5905 0.7546 0.7618 0.4471 0.6898 0.2780 0.5116 0.3571 0.5995 0.5544
22 (Ours) 0.5612 0.7530 0.7504 0.4573 0.6677 0.2880 0.5088 0.1001 0.5564 0.5159
21 (Ours) 0.5072 0.7470 0.7256 0.4019 0.6409 0.2840 0.4947 0.0311 0.4904 0.4803
20 (Ours) 0.4070 0.7410 0.7075 0.3942 0.6275 0.2900 0.4755 0.0311 0.4448 0.4576
Qwen3 4B (Instruct)
36 (Base) 0.7016 0.7644 0.8476 0.5879 0.6630 0.3720 0.5277 0.8499 0.4209 0.6372
31 (Ours) 0.6887 0.7459 0.8182 0.5367 0.6598 0.3520 0.5010 0.5375 0.3549 0.5772
30 (Ours) 0.6749 0.7497 0.8184 0.5239 0.6511 0.3340 0.4895 0.3677 0.3813 0.5545
29 (Ours) 0.6356 0.7470 0.7988 0.5043 0.6322 0.3300 0.4725 0.1797 0.3657 0.5184
28 (Ours) 0.5395 0.7432 0.7929 0.4898 0.6275 0.3000 0.4524 0.1266 0.3465 0.4909
27 (Ours) 0.4409 0.7296 0.7668 0.4454 0.5967 0.2960 0.4322 0.0356 0.2782 0.4468
26 (Ours) 0.3795 0.7062 0.7231 0.4002 0.5675 0.2840 0.4064 0.0318 0.3165 0.4239
Qwen3 14B (Instruct)
40 (Base ) 0.7883 0.8090 0.8775 0.6621 0.7451 0.4040 0.6133 0.8226 0.4652 0.6875
35 (Ours) 0.7792 0.7982 0.8695 0.6468 0.7356 0.3880 0.5825 0.6164 0.4376 0.6504
34 (Ours) 0.7742 0.7889 0.8561 0.6271 0.7388 0.3780 0.5792 0.7536 0.4700 0.6629
33 (Ours) 0.7588 0.7905 0.8561 0.6058 0.7269 0.3740 0.5693 0.6983 0.4376 0.6464
32 (Ours) 0.7370 0.7889 0.8481 0.5973 0.7103 0.3680 0.5575 0.6467 0.4017 0.6284
31 (Ours) 0.7156 0.7856 0.8552 0.5998 0.6859 0.3640 0.5412 0.4822 0.3801 0.6011
30 (Ours) 0.6543 0.7758 0.8384 0.5589 0.6732 0.3460 0.5197 0.2343 0.3657 0.5518
29 (Ours) 0.5954 0.7688 0.8215 0.5256 0.6575 0.3260 0.4945 0.1585 0.3381 0.5207
Table 2: Benchmark accuracy restoration through finetuning on Qwen3 4B with an effective depth of 27.
Finetuning Steps MMLU Arc C. GSM-8K (%)
0 (Ours) 0.4409 0.4454 0.0356
1024 (Ours) 0.6063 0.4932 0.4011
4096 (Ours) 0.6196 0.5102 0.4829
8192 (Ours) 0.6134 0.5162 0.4943
Qwen3 4B (Base) 0.7014 0.5887 0.8514
Figure 7: Wall clock time to complete the following inference tasks: KV Cache pre-filling, autoregressive generation, and single token generation with a pre-filled KV Cache. $\Delta$ indicates how many layers have been merged using LP (e.g. a $\Delta$ of 4 indicates that 2 groups of 2 layers have been converted to 2 effective layers). The gains in inference speed are directly proportional to the grade of LP. The 1-token generation task for Llama 3.2 3B does not saturate the GPU compute until a sequence length of 2048. Even in this regime, LP benefits from considerable speed-ups.
@@
Experimental protocol. For all our experiments, we use a node with two A100 SXM4 80Gb GPUs, four AMD EPYC 7742 CPUs, and 512Gb of RAM. We test for varying sequence lengths, up to $4096$ (Llama’s context window), with a batch size of $1$ unless indicated otherwise. We consider two models of the Llama family: Llama2 7B, and Llama3.2 3B, as well as two sizes from Qwen3: 4B and 14B. Given a desired effective depth, we replace the required number of normal layers with LP layers. For Llama 2 7B and 3.2 3B, the LP layers are selected based on the configuration that minimized the PPL for a given amount of LP (Fig. 6). For Qwen3, LP is applied until the 4th to last decoder layer. The rest of the layers implement the tensor parallel approach as described in [37]. For evaluation, we measure the ICL 5-shot accuracies using the lm-eval package [38]. We test the ICL accuracy of the models on several tasks: MMLU [39], PiQA [40], ARC Easy, ARC Challenge, Winogrande [41], OpenBookQA [42], Hellaswag [43], GSM-8K [44] and ifeval [45]. The perplexity of the models is always evaluated against a subset of the test set of RedPajama [1].
Impact of LP on PPL and ICL accuracies. We first examine how perplexity evolves when applying LP across layer sequences of different lengths and depths. Fig.6 reveals a common optimal sequence end-index minimizing perplexity, found at layers 28 for Llama2 7B and 25 for Llama3.2 3B. Table1 compares the In-Context Learning performance across models with varying effective depths. Performance declines gradually as LP increases, followed by a sharp drop beyond a certain threshold. Specifically, this occurs after reducing effective depth by 9 layers for Qwen3 14B, by 7 layers for Llama2 7B and Qwen3 4B, and by 5 layers for Llama3.2 3B. These results indicate that larger models are more robust to the computational graph modifications from LP, suggesting that our approach is likely applicable to current commercial-scale LLMs used in major deployments.
It is worth noting that, unlike the other benchmarks, GSM-8K already drops severely in performance when applying low amounts of LP. Recent mechanistic interpretability research shows that LLMs have special circuitry for math operations, localized in a small set of parameters [46, 47]. [48] identify math specific parameters, and report a drop in accuracy of 17% when pruning them. We believe that the changes in the computational graph by the use of LP in some of the late layers of the LLM interfere with these fragile and sparse subnetworks, while leaving general language competence largely intact.
@@
Fine-tuning for performance recovery. While LP provides speed improvements, associated architectural modifications may degrade model performance. To counteract this, we explored whether fine-tuning could effectively restore the original model’s capabilities. Specifically, we applied LP to layers 14–32 of Qwen3 4B, subsequently fine-tuning only these modified layers on randomly selected samples from the RedPajama training dataset [1]. Employing a batch size of 32, a linear learning rate schedule starting at $1e{-4}$, and the AdamW optimizer [49], we observed significant restoration of the model’s performance. Notably, GSM-8K accuracy improved from near-zero levels to more than half of the original model’s accuracy. It is possible that additional fine-tuning, or smarter tuning strategies, could yield further improvement, but resource constraints limited the scope of our experiments.
@@ 6 Limitations
The effectiveness of our approach exhibits notable variations across model scales. Smaller models show reduced benefits, likely due to their less sparse activation patterns and more tightly coupled layer dependencies. This degradation becomes more pronounced as the LP sequence increases, suggesting a practical upper limit to the number of layer pairs that can be effectively parallelized.
@@
Determining the ’true’ effective depth—the optimal configuration of parallel layer pairs—remains an open challenge as there is no theoretical framework for predicting the optimal grouping strategy.
These limitations highlight important directions for future research, particularly in developing more robust methods for determining optimal layer groupings and investigating the interplay between our approach and other efficiency-oriented techniques.